OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (3, -5) and is perpendicular to the line y = 3x - 6?

OpenStudy (anonymous):

If a line is perpendicular to another line, than the slops are opposite reciprocals or negative reciprocals. To find the equation you can use slope intercept form y-y1=m(x-x1), where m is the slope. In this case, the slope of the equation you want to find is -1/3 because that's the opposite reciprocal of 3, then just plug the point and slope into the formula. y-(-5)=-1/3(x-3) y+5=-1/3x+1 y=-1/3x-4